# Particle Semi-Implicit Variational Inference

### v1

@@ Appendix E Proofs in Section 3
Recall from Santambrogio [2015, Definition 7.12],
@@
for any perturbation $\chi=\tilde{p}-p$ with $\tilde{p}\in\mathcal{P}(\mathbb{R}^{d_{z}})\cap L_{c}^{\infty}(\mathbb{R}^{d_{z}})$ (see Santambrogio [2015, Notation]).
@@
where $\mathcal{E}^{\gamma}:(\theta,r)\mapsto\int\log\left(\frac{q_{\theta,r}(x)+\gamma}{p(x,y)}\right)q_{\theta,r}(\mathrm{d}x)$. Since $\delta_{r}\mathsf{R}_{\lambda}^{\mathrm{E}}[\theta,r]=\lambda_{r}\delta_{r}\mathrm{KL}(r|p_{0})$, its first variation follows immediately from standard calculations [Ambrosio et al., 2005, Santambrogio, 2015]. As for $\delta_{r}\mathcal{E}^{\gamma}$, we have the following proposition:
@@
$=$ $\int_{X}[q_{\theta,r}+\epsilon q_{\theta,\chi}](x)\log([q_{\theta,r}+\epsilon q_{\theta,\chi}](x)+\gamma)\,\mathrm{d}x$
$-\int_{X}[q_{\theta,r}+\epsilon q_{\theta,\chi}](x)\log{p(y,x)}\,\mathrm{d}x.$
@@
=$\overset{(a)}{=}\int_{Z}\left(\int_{X}k{{}_{\theta}}(x|z)\left[\log\left(\frac{q_{\theta,r}(x)+\gamma}{p(y,x)}\right)+\frac{q_{\theta,r}(x)}{q_{\theta,r}(x)+\gamma}\right]\,\mathrm{d}x\right)\chi(z)\,\mathrm{d}z.$
@@ Appendix F Proofs in Section 4
We first begin by proving $\Gamma$-convergence from the definition, i.e., demonstrating that the liminf inequality holds and establishing the existence of a recovery sequence. The latter follows from pointwise convergence:
@@
To arrive at the convergence of minima, we invoke Dal Maso [2012, Theorem 7.8] by using the fact that $\mathcal{E}_{\lambda}^{\gamma}$ is equi-coercive in the sense of Dal Maso [2012, Definition 7.6]. To see that $\mathcal{E}_{\lambda}^{\gamma}$ is equi-coercive, note that we have $\mathcal{E}_{\lambda}^{\gamma}\geq\mathcal{E}_{\lambda}$ and $\mathcal{E}_{\lambda}$ is l.s.c. (from Prop. 3), then applying Dal Maso [2012, Proposition 7.7]. ∎
@@
In Section F.4, we show that under our assumptions the drift $\tilde{b}^{\gamma}$ is Lipschitz. And under Lipschitz regularity conditions, the proof follows similarly to Lim et al. [2023] which we shall outline for completeness.
@@
whose existence and uniqueness of strong solutions are given by Carmona [2016][Thereom 1.2].
@@
$\mathsf{d}^{2}(F^{k}_{T}(\Upsilon)_{t},F^{k}_{T}(\Upsilon)_{t})\leq\frac{(tC(t))^{k}}{k!}\sup_{s\in[0,T]}\mathsf{d}^{2}(\Upsilon_{s},\Upsilon_{s}).$
@@
$\sup_{s\in[0,T]}\mathsf{d}^{2}(F^{k}_{T}(\Upsilon)_{s},F^{k}_{T}(\Upsilon)_{s})\leq\frac{(TC(T))^{k}}{k!}\sup_{s\in[0,T]}\mathsf{d}^{2}(\Upsilon_{s},\Upsilon^{\prime}_{s}).$
@@
where $\hat{r}^{\gamma}_{s}=\frac{1}{M}\sum_{m=1}^{M}\delta_{Z^{\gamma}_{s,m}}$ and we use Fournier and Guillin [2015]. Note that we also have
@@
Applying Eq. 26 in Eq. 25 then Eqs. 27 and 28, we obtain
@@
where $C_{z}:=3K_{b^{\gamma}}^{2}$and, as before, we apply Cauchy–Schwarz, Lipschitz and $C_{r}$ inequalities. Then from Eqs. 26, 27 and 28, we obtain
$(b)\leq C\mathbb{E}\int_{0}^{T}\sup_{s\in[0,T]}\|\theta^{\gamma,M}_{s}-\theta^{\gamma}_{s}\|^{2}+\sup_{s\in[0,T]}\frac{1}{M}\sum_{m=1}^{M}\|Z^{\gamma}_{s,m}-Z^{\gamma,M}_{s,m}\|^{2}\,+o(1)\mathrm{d}s,$ (30)
where $C:=C_{z}+2C_{\theta}$. Combining Eqs. 29 and 30 and applying Grönwall’s inequality, we obtain
@@
It is established below in Prop. 13 that $b$ satisfies a Lipschitz inequality, i.e., there is some $K_{b}\in\mathbb{R}_{>0}$ such that
@@
Under the same assumptions as Prop. 8, the function $(\theta,r)\mapsto\nabla_{\theta}\mathcal{E}_{\lambda}(\theta,r)$ is Lipschitz, i.e., there exist some constant $K_{\mathcal{E}_{\lambda}}\in\mathbb{R}_{>0}$ such that
@@
As we shall later show, that the two terms have the following upper bounds:
@@
which holds from the fact that $p_{k}$ has finite second moments Assumption 3, and the fact that $\mathbb{E}_{p_{k}(\epsilon)}\left\|d_{\theta,r}^{p,\gamma}(z^{\prime},\epsilon)\right\|$ is bounded.
@@
$\|b^{\gamma}(\theta,r,z)-b^{\gamma}(\theta^{\prime},r^{\prime},z^{\prime})\|$ $\leq\|\mathbb{E}_{p_{k}(\epsilon)}[(\nabla_{z}\phi_{\theta}\cdot[d_{\theta,r}^{p,\gamma}+\Gamma_{\theta,r}^{\gamma}])(z,\epsilon)-(\nabla_{z}\phi_{\theta^{\prime}}\cdot d_{\theta^{\prime},r^{\prime}}^{p}+\Gamma_{\theta^{\prime},r^{\prime}}^{\gamma})(z^{\prime},\epsilon)]\|$
@@
$\leq\mathbb{E}_{p_{k}(\epsilon)}\|(\nabla_{z}\phi_{\theta}\cdot d_{\theta,r}^{p,\gamma}+\Gamma_{\theta,r}^{\gamma})(z,\epsilon)-(\nabla_{z}\phi_{\theta^{\prime}}\cdot d_{\theta^{\prime},r^{\prime}}^{p,\gamma}+\Gamma_{\theta^{\prime},r^{\prime}}^{\gamma})(z^{\prime},\epsilon)\|$
@@
$\mathbb{E}_{p_{k}(\epsilon)}\|(\nabla_{z}\phi_{\theta}\cdot d_{\theta,r}^{p,\gamma}+\Gamma_{\theta,r}^{\gamma})(z,\epsilon)-(\nabla_{z}\phi_{\theta^{\prime}}\cdot d_{\theta^{\prime},r^{\prime}}^{p,\gamma}+\Gamma_{\theta^{\prime},r^{\prime}}^{\gamma})(z^{\prime},\epsilon)\|$
@@
$+$ $\mathbb{E}_{p_{k}(\epsilon)}\|\nabla_{z}\phi_{\theta}(z,\epsilon)\cdot[d_{\theta^{\prime},r^{\prime}}^{p,\gamma}+\Gamma_{\theta^{\prime},r^{\prime}}^{\gamma}](z^{\prime},\epsilon)-(\nabla_{z}\phi_{\theta^{\prime}}\cdot d_{\theta^{\prime},r^{\prime}}^{p,\gamma}+\Gamma_{\theta^{\prime},r^{\prime}}^{\gamma})(z^{\prime},\epsilon)\|$
@@
Under Assumption 2, the map $(\theta,r,z)\mapsto\Gamma^{\gamma}_{\theta,r}(z,\epsilon)$ is Lipschitz and bounded. (Lipschitz) there is constants $a_{\Gamma},b_{\Gamma}\in\mathbb{R}_{>0}$ such that following hold:
@@
This is immediate from the definition
@@
where (a) we add and subtract the relevant terms and apply Jensen’s inequality and the triangle inequality; (b) we use the fact that $\|\nabla_{x}q_{\theta^{\prime},r^{\prime}}\left(x^{\prime}\right)\|=\|\int\nabla_{x}k{{}_{\theta^{\prime}}}(x^{\prime}|z)r^{\prime}(\mathrm{d}z)\|\leq B_{k}$ (from Cauchy-Schwartz and the boundedness part of Assumption 2). Now, we deal with the terms individually. For the first term in Eq. 41, we use the fact that the map $(\theta,r,z)\mapsto\nabla_{x}q_{\theta,r}(\phi_{\theta}(z,\epsilon))$ is $K_{q}$-Lipschitz from Prop. 17. As for the second term in Eq. 41, we apply Prop. 18.
@@
$\overset{(b)}{\leq}K_{k}(\|(\theta,x)-(\theta^{\prime},x^{\prime})\|+\mathsf{W}_{1}(r,r^{\prime})),$
@@
$\overset{(a)}{\leq}B_{k}\mathsf{W}_{1}(r,r^{\prime})+B_{k}\|(\theta,x)-(\theta,x^{\prime})\|$
@@ G.1 Gradient estimator
$=\nabla_{z}\mathbb{E}_{\epsilon\sim p_{k}}\left[\log\frac{q_{\theta,r}(\phi_{\theta}(z,\epsilon))}{{p(y,\phi_{\theta}(z,\epsilon))}}+\frac{q_{\theta,r}(\phi_{\theta}(z,\epsilon))}{q_{\theta,r}(\phi_{\theta}(z,\epsilon))+\gamma}\right].$
@@
$\nabla_{z}\delta_{r}\mathcal{E}^{\gamma}[\theta,r](z)$ $=\mathbb{E}_{\epsilon\sim p_{k}}\left[\nabla_{z}\log\frac{q_{\theta,r}(\phi(z,\epsilon))}{{p(y,\phi(z,\epsilon))}}+\frac{\eta\nabla{q_{\theta,r}(\phi_{\theta}(z,\epsilon))}}{({q_{\theta,r}(\phi_{\theta}(z,\epsilon))+\gamma})^{2}}\right].$ (43)
@@ G.2 Preconditioners
$\mathrm{d}\theta_{t}=\Psi^{\theta}_{t}\nabla_{\theta}\mathcal{E}_{\lambda}(\theta_{t},r_{t}),\enskip\partial_{t}r_{t}=\nabla_{z}\cdot(r_{t}\Psi^{r}_{t}\nabla_{z}\delta_{r}\mathcal{E}_{\lambda}[\theta_{t},r_{t}]),$
@@
$=\nabla_{z}\cdot(r_{t}\Psi^{r}_{t}\nabla_{z}[\delta_{r}\mathcal{E}[\theta_{t},r_{t}]-\log p_{0}])+\nabla\cdot(r_{t}\Psi^{r}_{t}\nabla_{z}\log r_{t}).$
@@
$\nabla\cdot(r_{t}\Psi^{r}_{t}\nabla_{z}\log r_{t})$ $=\nabla_{z}\cdot(\Psi^{r}_{t}\nabla_{z}r_{t})=\nabla\cdot\left(\nabla\cdot[\Psi^{r}_{t}r_{t}]\right)-\nabla\cdot(r_{t}\nabla\cdot\Psi^{r}_{t})$
where $(\nabla\cdot\Psi^{r}_{t})_{i}=\sum_{j=1}^{d_{z}}\partial_{z_{j}}[(\Psi^{r}_{t})_{ij}]$. The last equality can be explicitly written as
@@
$\partial_{t}r_{t}=\nabla_{z}\cdot(r_{t}\left(\Psi^{r}_{t}\nabla_{z}[\delta_{r}\mathcal{E}[\theta_{t},r_{t}]-\log p_{0}]-\nabla\cdot\Psi^{r}_{t}\right))+\nabla\cdot\left(\nabla\cdot[\Psi^{r}_{t}r_{t}]\right))$
Examples. Following in the essence of RMSProp [Tieleman and Hinton, 2012], we utilize the preconditioner defined as follows:

### v3

@@ Appendix E Proofs in Section 3
Recall from Santambrogio, (2015, Definition 7.12),
@@
for any perturbation $\chi=\tilde{p}-p$ with $\tilde{p}\in\mathcal{P}(\mathbb{R}^{d_{z}})\cap L_{c}^{\infty}(\mathbb{R}^{d_{z}})$ (see Santambrogio, (2015, Notation)).
@@
where $\mathcal{E}^{\gamma}:(\theta,r)\mapsto\int\log\left(\frac{q_{\theta,r}(x)+\gamma}{p(x,y)}\right)q_{\theta,r}(\mathrm{d}x)$. Since $\delta_{r}\mathsf{R}_{\lambda}^{\mathrm{E}}[\theta,r]=\lambda_{r}\delta_{r}\mathrm{KL}(r|p_{0})$, its first variation follows immediately from standard calculations (Ambrosio et al.,, 2005; Santambrogio,, 2015). As for $\delta_{r}\mathcal{E}^{\gamma}$, we have the following proposition:
@@
$=$ $\int_{\mathcal{X}}[q_{\theta,r}+\epsilon q_{\theta,\chi}](x)\log([q_{\theta,r}+\epsilon q_{\theta,\chi}](x)+\gamma)\,\mathrm{d}x$
$-\int_{\mathcal{X}}[q_{\theta,r}+\epsilon q_{\theta,\chi}](x)\log{p(y,x)}\,\mathrm{d}x.$
@@
=$\overset{(a)}{=}\int_{\mathcal{Z}}\left(\int_{\mathcal{X}}k{{}_{\theta}}(x|z)\left[\log\left(\frac{q_{\theta,r}(x)+\gamma}{p(y,x)}\right)+\frac{q_{\theta,r}(x)}{q_{\theta,r}(x)+\gamma}\right]\,\mathrm{d}x\right)\chi(z)\,\mathrm{d}z.$
@@ Appendix F Proofs in Section 4
We first begin by proving $\Gamma$-convergence directly via its definition, i.e., demonstrating that the liminf inequality holds and establishing the existence of a recovery sequence. The latter follows from pointwise convergence:
@@
To arrive at the convergence of minima, we invoke Dal Maso, (2012, Theorem 7.8) by using the fact that $\mathcal{E}_{\lambda}^{\gamma}$ is equi-coercive in the sense of Dal Maso, (2012, Definition 7.6). To see that $\mathcal{E}_{\lambda}^{\gamma}$ is equi-coercive, note that we have $\mathcal{E}_{\lambda}^{\gamma}\geq\mathcal{E}_{\lambda}$ and $\mathcal{E}_{\lambda}$ is l.s.c. (from Prop. 3), then applying Dal Maso, (2012, Proposition 7.7). ∎
@@
In Section F.4, we show that under our assumptions the drift $\tilde{b}^{\gamma}$ is Lipschitz. And under Lipschitz regularity conditions, the proof follows similarly to Lim et al., (2024) which we shall outline for completeness.
@@
whose existence and uniqueness of strong solutions are given by Carmona, (2016)[Thereom 1.2].
@@
$\mathsf{d}^{2}(F^{k}_{T}(\Upsilon)_{t},F^{k}_{T}(\Upsilon^{\prime})_{t})\leq\frac{(tC(t))^{k}}{k!}\sup_{s\in[0,T]}\mathsf{d}^{2}(\Upsilon_{s},\Upsilon^{\prime}_{s}).$
@@
$\sup_{s\in[0,T]}\mathsf{d}^{2}(F^{k}_{T}(\Upsilon)_{s},F^{k}_{T}(\Upsilon^{\prime})_{s})\leq\frac{(TC(T))^{k}}{k!}\sup_{s\in[0,T]}\mathsf{d}^{2}(\Upsilon_{s},\Upsilon^{\prime}_{s}).$
@@
where $\hat{r}^{\gamma}_{s}=\frac{1}{M}\sum_{m=1}^{M}\delta_{Z^{\gamma}_{s,m}}$ and we use Fournier and Guillin, (2015). Note that we also have
@@
Applying Eq. 28 in Eq. 27 then Eqs. 29 and 30, we obtain
@@
where $C_{z}:=3K_{b^{\gamma}}^{2}$and, as before, we apply Cauchy–Schwarz, Lipschitz and $C_{r}$ inequalities. Then from Eqs. 28, 29 and 30, we obtain
$(b)\leq C\mathbb{E}\int_{0}^{T}\sup_{s\in[0,T]}\|\theta^{\gamma,M}_{s}-\theta^{\gamma}_{s}\|^{2}+\sup_{s\in[0,T]}\frac{1}{M}\sum_{m=1}^{M}\|Z^{\gamma}_{s,m}-Z^{\gamma,M}_{s,m}\|^{2}\,+o(1)\mathrm{d}s.$ (32)
Combining Eqs. 31 and 32 and applying Grönwall’s inequality, we obtain
@@
It is established below in Prop. 13 that $b^{\gamma}$ satisfies a Lipschitz inequality, i.e., there is some $K_{b^{\gamma}}\in\mathbb{R}_{>0}$ such that
@@
Under the same assumptions as Prop. 8, the function $(\theta,r)\mapsto\nabla_{\theta}\mathcal{E}^{\gamma}_{\lambda}(\theta,r)$ is Lipschitz, i.e., there exist some constant $K_{\mathcal{E}^{\gamma}_{\lambda}}\in\mathbb{R}_{>0}$ such that
@@
As we shall later show, the two terms have the following upper bounds:
@@
which holds from the assumption that $p_{k}$ has finite second moments Assumption 3, and from our assumption that $\mathbb{E}_{p_{k}(\epsilon)}\left\|d_{\theta,r}^{p,\gamma}(z^{\prime},\epsilon)\right\|$ is bounded. Hence, the map is Lipschitz and so LABEL:{eq:tlip_t1_b} holds.
@@
$\|b^{\gamma}(\theta,r,z)-b^{\gamma}(\theta^{\prime},r^{\prime},z^{\prime})\|$ $\leq\|\mathbb{E}_{p_{k}(\epsilon)}[(\nabla_{z}\phi_{\theta}\cdot[d_{\theta,r}^{p,\gamma}+\Gamma_{\theta,r}^{\gamma}])(z,\epsilon)-(\nabla_{z}\phi_{\theta^{\prime}}\cdot[d_{\theta^{\prime},r^{\prime}}^{p}+\Gamma_{\theta^{\prime},r^{\prime}}^{\gamma}])(z^{\prime},\epsilon)]\|$
@@
$\leq\mathbb{E}_{p_{k}(\epsilon)}\|(\nabla_{z}\phi_{\theta}\cdot d_{\theta,r}^{p,\gamma}+\Gamma_{\theta,r}^{\gamma})(z,\epsilon)-(\nabla_{z}\phi_{\theta^{\prime}}\cdot[d_{\theta^{\prime},r^{\prime}}^{p,\gamma}+\Gamma_{\theta^{\prime},r^{\prime}}^{\gamma}])(z^{\prime},\epsilon)\|$
@@
$\mathbb{E}_{p_{k}(\epsilon)}\|(\nabla_{z}\phi_{\theta}\cdot[d_{\theta,r}^{p,\gamma}+\Gamma_{\theta,r}^{\gamma})](z,\epsilon)-(\nabla_{z}\phi_{\theta^{\prime}}\cdot[d_{\theta^{\prime},r^{\prime}}^{p,\gamma}+\Gamma_{\theta^{\prime},r^{\prime}}^{\gamma}])(z^{\prime},\epsilon)\|$
@@
$+$ $\mathbb{E}_{p_{k}(\epsilon)}\|\nabla_{z}\phi_{\theta}(z,\epsilon)\cdot[d_{\theta^{\prime},r^{\prime}}^{p,\gamma}+\Gamma_{\theta^{\prime},r^{\prime}}^{\gamma}](z^{\prime},\epsilon)-(\nabla_{z}\phi_{\theta^{\prime}}\cdot[d_{\theta^{\prime},r^{\prime}}^{p,\gamma}+\Gamma_{\theta^{\prime},r^{\prime}}^{\gamma}])(z^{\prime},\epsilon)\|$
@@
Under Assumption 2, the map $(\theta,r,z)\mapsto\Gamma^{\gamma}_{\theta,r}(z,\epsilon)$ is Lipschitz and bounded. (Lipschitz) there are constants $a_{\Gamma},b_{\Gamma}\in\mathbb{R}_{>0}$ such that following hold:
@@
Let $x:=\phi_{\theta}(z,\epsilon)$, and $x^{\prime}:=\phi_{\theta^{\prime}}(z^{\prime},\epsilon)$. Then, we have
@@
where (a) we add and subtract the relevant terms and the triangle inequality; (b) we use the fact that $\|\nabla_{x}q_{\theta^{\prime},r^{\prime}}\left(x^{\prime}\right)\|=\|\int\nabla_{x}k{{}_{\theta^{\prime}}}(x^{\prime}|z)r^{\prime}(\mathrm{d}z)\|\leq B_{k}$ (from Cauchy–Schwartz and the boundedness part of Assumption 2). Now, we deal with the terms individually. For the first term in Eq. 43, we use the fact that the map $(\theta,r,z)\mapsto\nabla_{x}q_{\theta,r}(\phi_{\theta}(z,\epsilon))$ is $K_{q}$-Lipschitz from Prop. 17. As for the second term in Eq. 43, we apply Prop. 18.
@@
$\overset{(b)}{\leq}K_{gq}(\|(\theta,x)-(\theta^{\prime},x^{\prime})\|+\mathsf{W}_{1}(r,r^{\prime})),$
@@
$\overset{(a)}{\leq}K_{q}(\mathsf{W}_{1}(r,r^{\prime})+\|(\theta,x)-(\theta,x^{\prime})\|)$
@@ G.1 Gradient estimator
$=\nabla_{z}\mathbb{E}_{\epsilon\sim p_{k}}\left[\log\frac{q_{\theta,r}(\phi_{\theta}(z,\epsilon))+\gamma}{{p(y,\phi_{\theta}(z,\epsilon))}}+\frac{q_{\theta,r}(\phi_{\theta}(z,\epsilon))}{q_{\theta,r}(\phi_{\theta}(z,\epsilon))+\gamma}\right].$
@@
$\nabla_{z}\delta_{r}\mathcal{E}^{\gamma}[\theta,r](z)$ $=\mathbb{E}_{\epsilon\sim p_{k}}\left[\nabla_{z}\log\frac{q_{\theta,r}(\phi(z,\epsilon))+\gamma}{{p(y,\phi(z,\epsilon))}}+\frac{\gamma\nabla{q_{\theta,r}(\phi_{\theta}(z,\epsilon))}}{({q_{\theta,r}(\phi_{\theta}(z,\epsilon))+\gamma})^{2}}\right].$ (45)
@@ G.2 Preconditioners
$\mathrm{d}\theta_{t}=-\Psi^{\theta}_{t}\nabla_{\theta}\mathcal{E}_{\lambda}(\theta_{t},r_{t})\,\mathrm{d}t,\enskip\partial_{t}r_{t}=\nabla_{z}\cdot(r_{t}\Psi^{r}_{t}\nabla_{z}\delta_{r}\mathcal{E}_{\lambda}[\theta_{t},r_{t}]),$
@@
$=\nabla_{z}\cdot(r_{t}\Psi^{r}_{t}\nabla_{z}[\delta_{r}\mathcal{E}[\theta_{t},r_{t}]-\log p_{0}])+\nabla_{z}\cdot(r_{t}\Psi^{r}_{t}\nabla_{z}\log r_{t}).$
@@
$\nabla_{z}\cdot(r_{t}\Psi^{r}_{t}\nabla_{z}\log r_{t})$ $=\nabla_{z}\cdot(\Psi^{r}_{t}\nabla_{z}r_{t})=\nabla_{z}\cdot\left(\nabla_{z}\cdot[\Psi^{r}_{t}r_{t}]\right)-\nabla_{z}\cdot(r_{t}\nabla_{z}\cdot\Psi^{r}_{t})$
where $(\nabla\cdot\Psi^{r}_{t})_{i}=\sum_{j=1}^{d_{z}}\partial_{z_{j}}[(\Psi^{r}_{t})_{ij}]$, and last equality holds since
@@
$\partial_{t}r_{t}=\nabla_{z}\cdot(r_{t}\left(\Psi^{r}_{t}\nabla_{z}[\delta_{r}\mathcal{E}[\theta_{t},r_{t}]-\log p_{0}]-\nabla_{z}\cdot\Psi^{r}_{t}\right))+\nabla_{z}\cdot\left(\nabla_{z}\cdot[\Psi^{r}_{t}r_{t}]\right))$
Examples. Following in the essence of RMSProp (Tieleman and Hinton,, 2012), we utilize the preconditioner defined as follows: